OpenStudy (anonymous):

A wholesaler is offering two different packages of custom printed T-shirts and sweatshirts to groups for fund-raisers. One package contains 10 dozen T-shirts and 14 dozen sweatshirts for $1544. The other package contains 18 dozen T-shirts and 8 dozen sweatshirts for $1472. Write and solve a system of linear equations to find the cost, in dollars, of a dozen T-shirts. a. $76 c. $48 b. $58 d. $36

OpenStudy (imstuck):

Oooh, these are fun! We need two equations, one will represent the number of t-shirts and sweatshirts as a quantity, and the other will represent them as dollars. So here we go...

OpenStudy (imstuck):

Actually its not even that difficult. Here...

OpenStudy (imstuck):

Let's deal with package A first. It tells you there are 10 dozen t's. How many t's is that?

OpenStudy (anonymous):

120

OpenStudy (imstuck):

Good. Now do the same with the sweats. How many sweats is 14 dozen?

OpenStudy (anonymous):

168

OpenStudy (imstuck):

Yes. Good. Now that's out of the way. The thing here is that within package A you have 120 t's and 168 sweats that add up to $1544. The equation for that then is (I have t's as "t" and sweats as "s", ok?) 120t + 168s = 1544. That is equation one. Get that?

OpenStudy (anonymous):

yes

OpenStudy (imstuck):

Now let's do the second equation in the exact same way. You have 18 dozen t's in that package and 8 dozen sweats. How many t's and how many sweats, then?

OpenStudy (anonymous):

216 & 96

OpenStudy (imstuck):

Good! Now we have a system of equations. Keep in mind here that we have a cost PER t and a cost PER sweat, and in the end they want the cost for a DOZEN. So we will do this first and then convert at the end. No worries. Do you know how to solve systems of equations?

OpenStudy (imstuck):

Either the substitution method or the addition method will work here. I prefer substitution. Do you have a preference?

OpenStudy (anonymous):

sub. never did the addition method

OpenStudy (imstuck):

Ok, then good. I prefer the sub myself. Your system is as follows: 120t + 140s = 1544 216t + 96s = 1472. Which do you want to solve for first? The t's ok?

OpenStudy (anonymous):

yes

OpenStudy (imstuck):

Ok, this is a long and hairy process because it doesn't come out evenly but t = 417/360.

OpenStudy (imstuck):

Give me a sec...I have to run outside for a minute be right back!!!! Stay with me...ok?

OpenStudy (anonymous):

sure thing

OpenStudy (imstuck):

Ok, I'm back now.

OpenStudy (imstuck):

Let me check my math before i go on. I just want to make sure I'm correct.

OpenStudy (anonymous):

ok

OpenStudy (imstuck):

Actually, let's scrap that idea and go with the dozens of shirts. The numbers are smaller that way. Here's our new system then:

OpenStudy (imstuck):

10t + 14s = 1544 18t + 8s = 1472

OpenStudy (anonymous):

yes

OpenStudy (imstuck):

Good. Now here...doing the math and solving for t in the first equation gives you 10t = 1544 - 14s...and\[t=\frac{ 1544-14s }{ 10}\]That's what we will use for substitution into the second equation. It will be tricky because of the fraction. But oh well.

OpenStudy (anonymous):

so then what

OpenStudy (imstuck):

Subbing t into the second equation gives you this mess:\[18(\frac{ 1544-14s }{ 10 })+8s=1472\]Now you have to distribute the 18 into that:

OpenStudy (imstuck):

The trick here is to have a common denominator throughout the whole entire equation; that means that each term will have to be set to have a denominator of 10.

OpenStudy (imstuck):

\[\frac{ 27792 }{ 10 }-\frac{ 252s }{ 10 }+\frac{ 80s }{ 10 }=\frac{ 14720 }{ 10 }\]Believe it or not it will be easier to get the math done this way.

OpenStudy (anonymous):

im following continue

OpenStudy (imstuck):

Add your "s" terms and move the 27792/10 to the other side:\[-\frac{ 13072 }{ 10 }=-\frac{ 172s }{ 10 }\]Now you can cross multiply.

OpenStudy (imstuck):

When you do that you get -130720 = -1720s. Solving for s you get s = 76.

OpenStudy (imstuck):

That represents the cost of a dozen sweatshirts.

OpenStudy (anonymous):

Thank you very much IMStuck

OpenStudy (imstuck):

But you are asked to find the cost of a dozen t-shirts, so hold on.

OpenStudy (anonymous):

you really need to learn the addition method...also called the elimination method. Sometimes it is easier then dealing with fractions

OpenStudy (anonymous):

thanks for the tip

OpenStudy (imstuck):

10t + 14(76) = 1544...10t + 1064 = 1544...10t = 1544 - 1064...10t = 480...t = 48. So it costs $48 for a dozen t-shirts. That's choice c, even though it looks to me like you have a., c., b., d. Out of order? Anyway your answer is $48.

OpenStudy (imstuck):

kelliegirl33...he said he doesn't know the addition method. If you read the whole conversation first.

OpenStudy (imstuck):

Where he says, "sub. never did the addition method" from above. Yes it would be easier, but not if you don't know it.

OpenStudy (anonymous):

I did read the whole conversation. That is what we are here for...to teach. If he would have wanted to learn the addition method, I would have obliged. I meant no disrespect.

OpenStudy (imstuck):

I understand and did not mean to be short with you. I am a high school math teacher and if I am teaching my students to learn one method before another, there is a reason for it. And no...MOST of us here at this site are not teachers...we are only helpers.

OpenStudy (anonymous):

I understand your position and did not mean to step on your toes. I am sorry

OpenStudy (imstuck):

Np. Didn't mean anything at all. If you are here to help others understand, then you are doing an awesome thing!